# Daniel Calegari and Andrea Delgado

### v1

@@ 5 Threats to validity
In this section, we present the threats to validity according to 49: construction, internal, conclusion, and external.
As a threat to construct validity, the selection of sources or the search string may not retrieve all relevant articles. To minimize them, we selected recognized sources based on their coverage of international publications from conferences and top-level journals in the area, and we opted for a search string that maximized the inclusion of articles using process mining and education* as critical terms. Exploratory searches were carried out in the sources and discussion among the authors for prior adjustments.
Internal validity is threatened by reproducibility and bias in the selection of studies. To minimize them, the protocol was specified, including the sources, the search string, and the inclusion/exclusion criteria, and the results were discussed between the authors at each stage to prevent possible selection bias.
As a threat to external validity, the quantity and relevance of the selected studies can affect the generalization of results. To minimize this threat, the articles were selected using the inclusion/exclusion criteria defined in the protocol and discussions between the authors. Although the number of articles selected compared to those recovered may seem low, the review’s focus is specifically on EPM over curricula analysis, which narrows the area. The results presented represent the topic’s current state to our knowledge. In any case, there could be articles that need to be included that were not published or indexed at the time of the searches.
The validity of the conclusions is threatened by dependence on the point of view of the authors conducting the review. To minimize this threat, the information recovered from the analyzed articles is the one indicated in each proposal, and if there were different visions in the extraction of the identified categories, these were discussed between the authors.
@@ 6 Conclusions
We have summarized the results of a systematic literature review that identifies works on applying PM to curricular analysis. It has been a small but active research area in the last ten years, encompassing the PM community’s development.
Works span the three types of PM (discovery, conformance checking, and extension) and can be classified into five categories, from the discovery of educational trajectories to specific analysis of common problems such as stopout and dropout and the production of recommendations based on evidence.
The objectives cannot be reached using PM in isolation. Many other data mining and Artificial Intelligence (AI) techniques could be used in combination, such as clustering of traces for different student groups to perform more specific analysis, the definition of KPIs for expressing target educational goals, and using AI strategies to derive rules as recommendations for higher success rate study paths.
Despite the existing work, we have identified some research opportunities, e.g., tightening the connecting with EDM for improving curricular analysis, improving the methodological settings to perform a curricular analysis according to predefined questions of interest, and the development of openly-available tools for end users to reduce the costs of implementing a PM initiative from scratch.

### v2

@@ 5 Threats to validity
This section presents the threats to validity 56: construction, internal, conclusion, and external.
As a threat to construct validity, the selection of sources or the search string may not retrieve all relevant articles. To minimize them, we selected recognized sources based on their coverage of international publications from conferences and top-level journals in the area, and we opted for a search string that maximized the inclusion of articles using “process mining and education*” as critical terms. Exploratory searches were carried out in the sources and discussion among the authors for prior adjustments.
Reproducibility and bias in the selection of studies threaten internal validity. To minimize these, the protocol was specified, including the sources, the search string, and the inclusion/exclusion criteria, and the results were discussed between the authors at each stage to prevent possible selection bias.
As a threat to external validity, the quantity and relevance of the selected studies can affect the generalization of results. To minimize this threat, the articles were selected using the inclusion/exclusion criteria defined in the protocol and discussions between the authors. Although the number of chosen articles compared to those recovered may seem low, the review’s focus is specifically on EPM over curricula analysis, which narrows the area. The results presented represent the topic’s current state to our knowledge. There could be articles that need to be included that were not published or indexed at the time of the searches.
Dependence on the authors’ perspective of the review threatens the validity of the conclusions. To minimize this threat, the information recovered from the analyzed articles is the one indicated in each proposal. If different visions were used to extract the identified categories, these were discussed between the authors.
@@ 6 Conclusions
We have summarized the results of a systematic literature review that identifies works on applying PM to curricular analysis. It has been a small but active research area in the last ten years, encompassing the PM community’s development. The reviewed studies encompass the main types of PM—discovery, conformance checking, enhancement, and predictive/prescriptive analysis—and have been organized into five key categories: from uncovering educational trajectories to identifying issues such as dropout and stopout and generating evidence-based recommendations.
In general, the objectives cannot be reached using PM in isolation. Many other data mining and AI techniques could be used in combination, such as clustering of traces for different student groups to perform more specific analysis, the definition of KPIs to express target educational goals, and using AI to derive rules as recommendations for higher success rate study paths.
Despite progress, several research opportunities remain. Strengthening connections with Educational Data Mining (EDM) could enhance the depth of curricular analysis. Given the sensitive nature of student data, ensuring data privacy and addressing ethical concerns are critical considerations. Developing frameworks for secure data management and ethical guidelines would support the responsible use of information while respecting student privacy. Improvements in methodological frameworks are also essential to structure curricular analyses around targeted research questions. Flexible and scalable PM models that can adapt to the diversity and complexity of educational trajectories are essential to accommodate various institutional contexts and student behaviors. Another promising area is predictive modeling for early intervention, enabling institutions to identify at-risk students and take timely action based on early dropout or academic struggle indicators.
Additionally, developing openly accessible tools for non-technical users could lower implementation barriers, making PM-based insights more widely applicable within educational institutions. Real-time visualization tools and refined bottleneck analysis techniques could further support educational administrators in identifying and mitigating delays or inefficiencies within curricula. Cross-institutional comparisons, facilitated by standardized methodologies and metrics, would enable universities to benchmark and enhance their curricular practices, providing a broader perspective on educational effectiveness and impact. Lastly, detecting concept drift—or significant changes in student progression patterns over time—could help institutions continuously adapt their curricula to evolving educational needs and student demographics.
Funding Information
No funding has been received for this review.
Conflict of Interest Statement
The authors have declared no conflicts of interest for this article.
Data Availability Statement
The list of selected studies is provided as supplementary material at http://dx.doi.org/10.5281/zenodo.14253621
ORCID